OpenStudy (anonymous):

Solve. Please show all of your work. During one year, the Hills' real estate bill included $413 for miscellaneous services. Of this amount, 8% went to the local homeless shelter. How much money did the homeless shelter receive?

OpenStudy (bahrom7893):

413-100% x - 8% x = 8*413/100 = 33.04

OpenStudy (anonymous):

413 * .08 my friend.

OpenStudy (anonymous):

awesome